OpenStudy (anonymous):

Rationalize the denominator and simplify. A. 34 B. 9 C. D.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

c\[81\sqrt{5\div5}\]

OpenStudy (anonymous):

5 divide by 5 will be..?

OpenStudy (anonymous):

9\[9\sqrt{5}\div5\]

OpenStudy (anonymous):

they are suppose to be on the bottom of 5

OpenStudy (anonymous):

i think its d.

OpenStudy (anonymous):

dang it the question didnt come up.

OpenStudy (anonymous):

\[\sqrt{_{5}^{81}}\]